y(t) + 4 y(t) = 6 e^2t with y(0) = 3 find solution of this differential equation with use laplace transform.
1
Expert's answer
2020-04-21T13:36:13-0400

In order to solve differential equation using Laplace transform, one has to take Laplace transform of both sides of equation, using the properties of the transform, solve for image "Y(s) \equiv L[y(t)]", and take inverse Laplace transform.

Using property "L[y'(t)] = s Y(s) - y(0)", table Laplace transform "L(e^{a t}) = \frac{1}{s-a}", and taking transform of both sides of "y'(t) + 4 y(t) = 6 e^{2 t}", obtain:

"s Y(s) - y(0) + 4 Y(s) = \frac{6}{s-2}".

Using initial condition "y(0) = 3" and simplifying, obtain "Y(s)[s + 4] = \frac{6}{s-2} + 3 = \frac{3 s}{s-2}", from where "Y(s) = \frac{3 s}{(s+4)(s-2)}".

In order to take inverse Laplace transform, one has to expand the last expression into partial fractions: "\frac{3 s}{(s+4)(s-2)} = \frac{A}{s+4} + \frac{B}{s-2}". Putting terms in the right hand side to common denominator, and equating the coefficients next to powers of "s" in the numerator of left and right hand side, obtain system of linear equations for "A, B":

"A+B = 3; -2 A + 4 B = 0", which has roots "A = 2; B = 1".

Hence, "\frac{3 s}{(s+4)(s-2)} = \frac{2}{s+4} + \frac{1}{s-2}".

From table transform "L[e^{a t}] = \frac{1}{s-a}", the inverse of expressions like "\frac{1}{s-a}" is "L^{-1}[\frac{1}{s-a}] = e^{a t}".

Therefore, the solution is "y(t) = L^{-1}[Y(s)] = L^{-1}[\frac{2}{s+4} + \frac{1}{s-2}] = 2 e^{-4 t} + e^{2 t}".